OpenStudy (anonymous):

help..... Mass box A = 10 grams; Mass box B = 5 grams; Mass box C—made of one A and one B How much A and B would be needed to produce 30 grams of C?_____ g of A; _______ g of B.

OpenStudy (anonymous):

we have a Proportion now !;) We have : A_________10______x B_________5_______y C_________15_____30 we have : 15*2=30 So have *2 for all .

OpenStudy (anonymous):

have : 10*2=A 5*2=B Got it ? @khofer

OpenStudy (anonymous):

I hope can help you ;)

OpenStudy (anonymous):

you did thank u:).....